### v2

# Hunting for a 17 MeV particle coupled to electrons

Abstract
We discuss a set of precision observables that can probe the existence of a light particle $X$ coupled to electrons in the mass range of 1–100 MeV. As a case study, we consider the recent excess of $e^{+}e^{-}$ final-state events at $\sqrt{s}=16.9$ MeV reported by the PADME collaboration. Interestingly, this mass is tantalizingly close to the invariant mass at which anomalous $e^{+}e^{-}$ pair production has previously been observed in nuclear transitions from excited to ground states by the ATOMKI collaboration. For the scenario in which the new particle has a vector coupling to electrons, we show that the PADME excess is already in tension with constraints from the anomalous magnetic moment of the electron and the non-observation of the exotic decay $\pi^{+}\to e^{+}\nu X$ at the SINDRUM experiment. Further improvements in the measurement of the electron $g$-2, together with upcoming results from PIONEER (searching for $\pi^{+}\to e^{+}\nu X$) and Mu3e (searching for $\mu^{+}\to e^{+}\overline{\nu}_{\mu}\nu_{e}X$), are expected to definitively probe this scenario in the near future. We also explore alternative possibilities where the new particle has scalar, pseudoscalar, or axial-vector couplings.
I Introduction
The PADME collaboration has recently reported an excess of $e^{+}e^{-}$ events peaked at $\sqrt{s}=16.90$ MeV, with a moderate significance of $1.8\,\sigma$ (global) and $2.5\,\sigma$ (local) Bossi et al. (2025); T. Spadaro, on behalf of the PADME collaboration (2025a, b). This energy scale matches the invariant mass at which anomalous $e^{+}e^{-}$ pair production has previously been observed, with high statistical significance, in nuclear transitions from excited to ground states in 8Be Krasznahorkay et al. (2016, 2018), 4He Krasznahorkay et al. (2019a, b, 2021, 2023); Krasznahorky et al. (2024), and 12C Krasznahorkay et al. (2022) by the ATOMKI collaboration, as well as in an experiment conducted at the VNU University of Science involving 8Be Anh et al. (2024). Such a coincidence of scales significantly reinforces the case for a common underlying origin, potentially involving a hypothetical new boson, commonly referred to as $X_{17}$ (see e.g. Feng et al. (2016, 2017, 2020); Zhang and Miller (2021); Barducci and Toni (2023); Denton and Gehrlein (2023); Alves et al. (2023); Abdallah et al. (2024)).
The MEG II experiment at PSI Afanaciev et al. (2024) also searched for the $X_{17}$ particle using the same 7Li($p$, $e^{+}e^{-}$) 8Be reaction employed by ATOMKI. While no significant signal was observed, the results are compatible with ATOMKI within $1.5\,\sigma$, and can be consistently included in a global fit Barducci et al. (2025). Adding also the recent PADME result Bossi et al. (2025); T. Spadaro, on behalf of the PADME collaboration (2025a, b) one obtains the best-fit value $m_{X}=16.88(5)$ MeV Arias-Aragón et al. (2025).
Currently, there is no clear indication of the spin-parity nature of the $X_{17}$ particle. Scalar or pseudoscalar couplings to nucleons fail to simultaneously account for all observed nuclear transitions, while vector and axial-vector scenarios face tensions with other low-energy constraints on the $X_{17}$ couplings to nucleons (see e.g. Zhang and Miller (2021); Barducci and Toni (2023); Hostert and Pospelov (2023)). In contrast, the coupling to electrons, relevant only through the branching ratio of $X_{17}$ into $e^{+}e^{-}$, plays only a minor role in explaining the nuclear decay anomalies, but it can be directly tested by PADME Nardi et al. (2018); Darmé et al. (2022); Arias-Aragón et al. (2024); Bertelli et al. (2025).
While other beam-dump experiments in the $X_{17}$ mass range, such as NA64 Banerjee et al. (2020); Andreev et al. (2021), could directly probe the $X_{17}$ coupling to electrons, potentially closing the entire available parameter space Beacham et al. (2020), it remains important to further explore complementary and independent tests that could confirm or disprove the PADME excess.
In this work, we examine a set of observables sensitive to a new light particle coupled to electrons in the 1–100 MeV mass range. While exploring in more generality various coupling scenarios (scalar, pseudoscalar, vector, and axial-vector) we focus in particular on a 17 MeV state with vector couplings to electrons, motivated by the recent PADME analysis. Specifically, we assess current constraints and future sensitivities from the electron anomalous magnetic moment, exotic pion decays at the SINDRUM experiment Eichler et al. (1986), as well as upcoming results from PIONEER Altmannshofer et al. (2022) and Mu3e Arndt et al. (2021), which aim to search for exotic decays of the pion and muon, respectively. These experiments are expected to decisively test the above scenario in the near future.
The paper is structured as follows. In Sects. IIV, we introduce the relevant observables, providing a brief overview of the key formulae and the expected experimental sensitivities. Our results are presented in Sect. VI, along with a description of the underlying phenomenological analysis. We conclude in Sect. VII, summarizing the main findings and outlining the prospects for testing the $X_{17}$ hypothesis through its coupling to electrons.
II Phenomenological setup
Let us define the interaction Lagrangian of a new light state, $X$, with electrons as
$\mathcal{L}_{X}\supset g_{eX}\left(\overline{e}\,\Gamma_{X}e\right)X\,,$ (1)
where $X=S,P,V,A$, and $\Gamma_{S}=1$ if $S$ is a scalar, $\Gamma_{P}=i\gamma_{5}$ if $P$ is a pseudoscalar, $\Gamma_{V}=\gamma_{\mu}$ if $V_{\mu}$ is a vector, and $\Gamma_{A}=\gamma_{\mu}\gamma_{5}$ if $A_{\mu}$ is an axial-vector.
In the following, we consider several observables that can probe the interaction of the $X$ particle with electrons, as described in Eq. (1).
For a vector coupling, the best-fit value associated with the $e^{+}e^{-}$ excess observed by the PADME collaboration is given by Bossi et al. (2025); T. Spadaro, on behalf of the PADME collaboration (2025a, b):
$g_{eV}=5.6\times 10^{-4}\,,\quad m_{V}=16.9\ \text{MeV}\,.$ (2)
We will use this benchmark point to assess its compatibility with existing experimental constraints.
III Electron $\boldsymbol{g-2}$
The anomalous magnetic moment of the electron, $a_{e}\equiv(g_{e}-2)/2$, has been commonly used to extract the value of the fine-structure constant, $\alpha$. However, recent improvements in atomic-physics experiments using Cesium (Cs) and Rubidium (Rb) interferometry have led to the following results for $\alpha$:
[…]
$(\Delta a_{e})_{\rm Cs}$ $=(-102.0\pm 26.4)\times 10^{-14}\,,$ (5)
$(\Delta a_{e})_{\rm Rb}$ $=\,\,\,\,\,\,\,\,(33.8\pm 16.1)\times 10^{-14}\,.$ (6)
It is interesting to consider the dominant sources of error in the determinations of $(\Delta a_{e})_{\rm Cs}$ and $(\Delta a_{e})_{\rm Rb}$, which stem from the atomic measurement of $\alpha$ and the experimental measurement of $a^{\rm exp}_{e}$ Di Luzio et al. (2025):
$\Delta a_{e}$ error source Value       % of $\Delta a_{e}$ error
$\alpha({\rm Cs})$, $\delta a_{e}^{\alpha({\rm Cs})}$ $22\times 10^{-14}$ $70\%$
$\alpha({\rm Rb})$, $\delta a_{e}^{\alpha({\rm Rb})}$ $9\times 10^{-14}$ $28\%$
Experiment, $\delta a_{e}^{\rm exp}$ $13\times 10^{-14}$ $24\%\,({\rm Cs})/59\%\,({\rm Rb})$
An improvement on $\delta a_{e}^{\rm exp}$ by a factor of $\sim 5$ is expected in the next few years. [^1] G. Gabrielse, private communication. On the same timescale, new measurements of $\alpha({\rm Rb})$ and $\alpha({\rm Cs})$ will significantly reduce the systematic effects that were dominant in their previous measurements, hopefully resolving the current discrepancy. [^2] S. Guellati-Khélifa and H. Mueller, private communications.
Therefore, $a_{e}$ can be used as a precision test of the SM and its extensions Giudice et al. (2012); Crivellin et al. (2018); Erdelyi et al. (2025). The contribution of the boson $X$ to the electron $g$-2 stemming from the Lagrangian of Eq. (1) reads Jegerlehner (2017); Athron et al. (2021):
$\Delta a^{X}_{e}=\frac{g^{2}_{eX}}{4\pi^{2}}\frac{m^{2}_{e}}{m^{2}_{X}}\,L_{X}\,,$ (7)
where, in the limit $m_{X}\gg m_{e}$, it turns out that
$L_{S}$ $=\ln\frac{m_{S}}{m_{e}}-\frac{7}{12}\,,$ (8)
$L_{P}$ $=-\ln\frac{m_{P}}{m_{e}}+\frac{11}{12}\,,$ (9)
$L_{V}$ $=\frac{1}{3}\,,~{}~{}L_{A}=-\frac{5}{3}\,,$ (10)
leading to the following numerical estimates:
$\!\!\!\Delta a^{S}_{e}$ $\!\approx 2.1\!\times\!10^{-11}\!\left(\frac{g_{eS}}{5.6\!\times\!10^{-4}}\right)^{\!2}\!\!\left(\frac{17\,{\rm MeV}}{m_{S}}\right)^{\!2}\!,$ (11)
$\!\!\!\Delta a^{P}_{e}$ $\!\approx-1.9\!\times\!10^{-11}\!\left(\!\frac{g_{eP}}{5.6\!\times\!10^{-4}}\!\right)^{\!2}\!\!\left(\!\frac{17\,{\rm MeV}}{m_{P}}\!\right)^{\!2}\!,$ (12)
$\!\!\!\Delta a^{V}_{e}$ $\!\approx 2.4\!\times\!10^{-12}\!\left(\frac{g_{eV}}{5.6\!\times\!10^{-4}}\right)^{\!2}\!\!\left(\frac{17\,{\rm MeV}}{m_{V}}\right)^{\!2}\!,$ (13)
$\!\!\!\Delta a^{A}_{e}$ $\!\approx-1.2\!\times\!10^{-11}\!\left(\!\frac{g_{eA}}{5.6\!\times\!10^{-4}}\!\right)^{\!2}\!\!\left(\!\frac{17\,{\rm MeV}}{m_{A}}\!\right)^{\!2}\!\!,$ (14)
where the PADME best-fit value of Eq. (2), which refers only to the vectorial case, has been assumed also for the other scenarios just as a reference.
Comparing the results of Eqs. (11-14) with Eqs. (5) and (6) we learn that the PADME result, see Eq. (2), is in a significant tension with the electron $g$-2 bound. Hereafter, for definiteness, we assume three benchmark scenarios: 1) $|\Delta a_{e}|\leq 10^{-12}$, where we inflated the current experimental errors in $\delta a^{\alpha(Cs)}_{e}$ and $\delta a^{\alpha(Rb)}_{e}$ to make Eqs. (5) and (6) consistent, 2) $|\Delta a_{e}|\leq 10^{-13}$, assuming a resolution of the current discrepancy in the measurements of $\alpha$ with a precision of $\mathcal{O}(10^{-13})$, and 3) $|\Delta a_{e}|\leq 10^{-14}$, which is the ultimate expected uncertainty on $\Delta a_{e}$ if both $\delta a^{\rm exp}_{e}$ and $\delta a^{\alpha}_{e}$ will improve by roughly one order of magnitude Di Luzio et al. (2025).
IV Pion decay
Another way to detect the $X$ boson is through the charged pion decay $\pi^{+}\to e^{+}\nu X$ [^3] For a related analysis, including also vector and axial-vector couplings to quarks, see Ref. Hostert and Pospelov (2023). The main focus here is on the electron coupling and the interplay of this observable with the PADME excess. The SINDRUM experiment searched for $e^{+}e^{-}$ resonances in $\pi^{+}\to e^{+}\nu X$, $X\to e^{+}e^{-}$ with sensitivity to branching ratios of $\mathcal{O}(10^{-10})$ Eichler et al. (1986). In order to constrain the parameter space of our scenarios, we exploit the bounds by SINDRUM Eichler et al. (1986) given as a function of the $X$ mass and lifetime. If $X$ is invisible, a complementary limit can be derived from a search for $\pi^{+}\to e^{+}\nu X$ at PIENU Aguilar-Arevalo et al. (2021). However, in the parameter space of our interest, we find that $X$ decays promptly into the visible $e^{+}e^{-}$ channel and therefore the PIENU bounds do not apply in our case. In the future, charged pion decays into visible final states can be searched for at the proposed experiment PIONEER Altmannshofer et al. (2022), planning to reach a sensitivity to branching ratios of $\mathcal{O}(10^{-11})$.
To assess the expected sensitivity on the coupling $g_{eX}$, it is useful to estimate the decay rate of $\pi^{+}\to e^{+}\nu X$ in naive dimensional analysis (NDA), assuming $m_{X}\ll m_{\pi^{+}}$.
When $X=S,P$, we find that $\Gamma(\pi^{+}\to\mu^{+}\nu X)\sim g^{2}_{eX}G^{2}_{F}m^{3}_{\pi}f^{2}_{\pi}|V_{ud}|^{2}/(4\pi)^{3}$, where $G_{F}$ is the Fermi constant, $f_{\pi}$ the pion decay constant, $V_{ud}$ the relevant CKM matrix element, and $1/(4\pi)^{3}$ stands for the three body phase space. Therefore, we find that
$\!\!\!\mathcal{B}(\pi^{+}\!\to\!e^{+}\nu X)\simeq\frac{\Gamma(\pi^{+}\!\to\!e^{+}\nu X)}{\Gamma(\pi^{+}\!\to\!\mu^{+}\nu)}\!\approx\frac{m^{2}_{\pi}}{m^{2}_{\mu}}\frac{g^{2}_{eX}}{16\pi^{2}}\,,$ (15)
where we used the standard pion decay rate $\Gamma(\pi^{+}\to\mu^{+}\nu)\approx G^{2}_{F}m_{\pi}m^{2}_{\mu}f^{2}_{\pi}|V_{ud}|^{2}/4\pi$. The current experimental bound $\mathcal{B}(\pi^{+}\!\to\!e^{+}\nu X)\lesssim 6\times 10^{-10}$ Eichler et al. (1986) by SINDRUM is therefore expected to translate into a bound on the coupling $g_{eX}\lesssim\,\text{few}\times 10^{-4}$.
On the other hand, if $X=V,A$ is a gauge boson coupled to a non-conserved current in an effective field theory where gauge invariance is restored at a scale $\Lambda$, its longitudinal mode can participate in processes at energies $E<\Lambda$, leading to enhancements by factors of $\mathcal{O}(E^{2}/m_{X}^{2})$. In the context of radiative pion decays, this implies an additional enhancement of $\mathcal{O}(m_{\pi}^{2}/m_{X}^{2})$, which places even tighter constraints on light spin-1 mediators Dror et al. (2017); Dror (2020); Hostert and Pospelov (2023). [^4] We note that in the case of a conserved current, such as for a dark photon, the enhancement discussed above is absent, and the resulting bounds from pion decay are expected to be comparatively weaker. Although such a scenario is not captured by the simplified model in Eq. (1), the introduction of additional couplings of $X$ to other SM fermions would require the inclusion of further observables.
The exact analytic expressions for the differential decay rates of leptonic meson decays accompanied by the emission of an axion-like particle have been derived in Refs. Aditya et al. (2012); Gallo et al. (2022); Altmannshofer et al. (2023). We have verified that MadGraph5_aMC@NLO Alwall et al. (2014) accurately reproduces these known results, and we have further utilized it to obtain the corresponding predictions for the scalar and (axial-)vector cases. For the representative benchmark of $m_{X}=17$ MeV, and normalising to the SINDRUM sensitivity $\mathcal{B}(\pi^{+}\to e^{+}\nu X)\lesssim 6\times 10^{-10}$ at this $X$ mass, we find that
$\!\!\frac{\mathcal{B}(\pi^{+}\to e^{+}\nu X)}{6\times 10^{-10}}$ $\approx 3.9\left(\frac{g_{eX}}{5.6\!\times\!10^{-4}}\right)^{\!2}\,,$ (16)
when $X=S,P$, in good agreement with NDA in Eq. (15). Similarly, when $X=V,A$, we find
$\!\!\frac{\mathcal{B}(\pi^{+}\to e^{+}\nu X)}{6\times 10^{-10}}$ $\approx 87\left(\frac{g_{eX}}{5.6\!\times\!10^{-4}}\right)^{\!2}\,,$ (17)
again in good agreement with NDA after taking into account the enhancement factor $\mathcal{O}(m_{\pi}^{2}/m_{X}^{2})$ discussed above.
V Muon decay
Muon decays could also be exploited to probe a light particle $X$ coupled to electrons Echenard et al. (2015); Knapen et al. (2024a, b). In particular, the upcoming Mu3e experiment at PSI searching for lepton flavour violation (LFV) in $\mu^{+}\to e^{+}e^{-}e^{+}$ will benefit from an intense muon beam, producing and stopping $\mathcal{O}(10^{15})$ muons during Phase I, with an order of magnitude improvement in Phase II Arndt et al. (2021).
The signature of interest in our analysis differs from the LFV process $\mu^{+}\to e^{+}e^{-}e^{+}$. Instead, we focus on the scenario in which a light particle $X$ is radiated from a final-state positron in the decay $\mu^{+}\to e^{+}\overline{\nu}_{\mu}\nu_{e}X$, with $X$ subsequently decaying into an $e^{+}e^{-}$ pair. As a result, the invariant mass of the electron and one of the positrons in the final state satisfies the relation $(p_{e^{+}}+p_{e^{-}})^{2}=m^{2}_{X}$. The experimental signature of this process is therefore characterized by a narrow bump in the invariant mass spectrum of $e^{+}e^{-}$ pairs, appearing on top of the SM background from the five-body decay $\mu^{+}\to e^{+}e^{-}e^{+}\overline{\nu}_{\mu}\nu_{e}$.
In addition to analyzing the invariant mass distributions of the $e^{+}e^{-}$ pairs for resonance searches, Ref. Knapen et al. (2024a) investigated the impact of incorporating angular observables. This inclusion was shown to enhance the expected experimental sensitivity to the (pseudo)scalar-electron coupling $g_{eX}$, with $X=S,P$, by a factor of $\mathcal{O}(1)$. For the case of such light states, we adopt the projected sensitivities from Ref. Knapen et al. (2024a) and present them in Fig. 1 (upper panel), using different shades of blue to indicate the corresponding Mu3e phase. In contrast, for the case of (pseudo)vector mediators, Ref. Knapen et al. (2024a) does not provide projections assuming couplings exclusively to electrons. To address this, we adopt the projected sensitivity to ${\rm Br}(\mu^{+}\to e^{+}\nu_{e}\overline{\nu}_{\mu}V,V\to e^{+}e^{-})$ from Ref. Perrevoort (2018) and compute the corresponding branching ratio using MadGraph5_aMC@NLO, considering only radiation from the positron. The resulting constraints on the coupling $g_{eX}$, with $X=V,A$ are shown in Fig. 1 (lower panel).
In comparison to pion decay experiments such as SINDRUM and PIONEER, the projected sensitivity of Mu3e is consistently weaker. This difference arises from the chirality suppression present in the SM background $\pi^{+}\to e^{+}\nu_{e}\gamma^{*},\gamma^{*}\to e^{+}e^{-}$, which does not apply to muon decays.
Additionally, we carried out an independent sensitivity study using the FeynRulesUfoMadGraph5_aMC@NLO toolchain Alloul et al. (2014); Degrande et al. (2012); Alwall et al. (2011, 2014) to simulate both signal and background processes for $\mu^{+}\to e^{+}e^{-}e^{+}\nu_{e}\overline{\nu}_{\mu}$. Adopting a fixed mass resolution of $0.5$ MeV across the relevant invariant mass range as in Knapen et al. (2024a), we perform a bin-by-bin analysis over the mediator mass range $m_{X}\in[2\,,100]$ MeV. For each bin, we apply Poisson statistics to derive upper bounds on the coupling $g_{eX}$. Given the large number of background events $B$ per bin, the Gaussian approximation is applicable, and we quantify sensitivity using $S/\sqrt{B}\simeq 1.64$, which corresponds approximately to a 95% confidence level (CL), under the assumption that statistical uncertainties dominate. We confirm the results of Perrevoort (2018); Knapen et al. (2017) up to $\mathcal{O}(1)$ differences attributable to detector effects.
VI Results
Figure 1: Present constraints (solid lines) and expected sensitivities (dashed lines) for a light $X=S,P,V,A$ particle coupled to electrons in the 1–100 MeV mass range. The vertical, red line at $m_{X}=16.9$ MeV denotes the $X_{17}$ benchmark, with the red star (in the vector case) representing the PADME best-fit value $g_{eV}=5.6\times 10^{-4}$.
The results of our analysis are displayed in Fig. 1, which shows the present constraints (solid lines) and expected sensitivities (dashed lines) for a light $X=S,P,V,A$ particle coupled to electrons in the 1–100 MeV mass range.
When the light particle is a spin-0 state, we find the limits to be comparable between the scalar ($X=S$ in the upper-left panel of Fig. 1) and the pseudoscalar ($X=P$ in the upper-right panel). In both cases, the most stringent constraints arise from measurements of the electron $g$-2 (dashed lines in different shades of green), with the precision in $|\Delta a_{e}|$ of $\mathcal{O}(10^{-14})$ allowing to probe the (pseudo)scalar couplings at the level of $g_{eX}\lesssim 10^{-5}$. A more moderate precision of $|\Delta a_{e}|\simeq\mathcal{O}(10^{-13})$ results in bounds which overlap to a good extent with the projections at the proposed PIONEER experiment which aims to reach sensitivities in radiative pion decays at the level of $\mathcal{B}(\pi^{+}\to e^{+}\nu X)\lesssim\mathcal{O}(10^{-11})$ (dashed orange line). Interestingly, the current bounds from the SINDRUM experiment (solid orange line) searching for the neutral particle emitted in $\pi^{+}$ decay approximately correspond to the sensitivity in $|\Delta a_{e}|\simeq\mathcal{O}(10^{-12})$ in the mass region above 40 MeV. At present, these constitute the most sensitive probes of new physics scenarios coupled to electrons in the 1–100 MeV mass range. Future muon experiments are expected to reach a sensitivity of $g_{eX}\lesssim 10^{-4}$ through the peak search in the invariant mass spectrum of $e^{+}e^{-}$ pair originating from a narrow resonance $X$, supplemented by the angular observables constructed from the $e^{\pm}$ momenta. The projected limits from the Mu3e experiment are shown as dashed blue lines, accounting for two planned phases: Phase I, with $\mathcal{O}(2.5\times 10^{15})$ muon decays, and Phase II, with an expected total of $\mathcal{O}(5.5\times 10^{16})$ events.
When the light particle is a spin-1 state, the limits are overall more stringent than in the spin-0 case. This is particularly evident for the constraints from exotic $\pi^{+}$ decays, as set by SINDRUM (solid orange) and projected PIONEER (dashed orange) experiments. These stronger bounds arise due to the additional enhancement from the emission of the longitudinal mode, which contributes a factor of $\mathcal{O}(m_{\pi}^{2}/m_{X}^{2})$ to $\mathcal{B}(\pi^{+}\to e^{+}\nu X)$. This enhancement is especially significant in the low-mass region when compared with the spin-0 scenarios. Indeed, these pion decay limits represent the most stringent constraints, followed by those from projections on the precision of $\Delta a_{e}$ (dashed green lines), reversing the pattern observed in the spin-0 case. A similar trend holds for experiments utilizing intense muon beams: in Phase II of the Mu3e experiment, the spin-1 scenario is expected to be probed down to $g_{eX}\simeq 8\times 10^{-5}$ at the benchmark mass of $m_{X}=17$ MeV, compared to $g_{eX}\simeq 10^{-4}$ in the spin-0 case. Furthermore, the axial-vector scenario ($X=A$, shown in the lower-right panel of Fig. 1) is more tightly constrained by the electron $g$-2 than the vector case ($X=V$, shown in the lower-left panel). Remarkably, the PADME best-fit value of $g_{eV}=5.6\times 10^{-4}$ (indicated by the red star in Fig. 1) is already in tension with these constraints, and upcoming results from Mu3e and PIONEER will be well positioned to robustly test this scenario.
VII Conclusions
In this work, we have tested the hypothesis of a new 17 MeV boson coupled to electrons, commonly referred to as $X_{17}$, by employing a set of observables complementary to those probed by PADME and other beam-dump experiments. We have shown that the $e^{+}e^{-}$ PADME excess is already in tension with the electron $g$-2, as well as exotic pion decay data from SINDRUM, although it remains marginally viable and testable.
We have analyzed a set of observables sensitive to the coupling of a light particle to electrons in the 1–100 MeV mass range, identifying regions of parameter space that are currently allowed and those soon to be probed. In particular, future measurements from Mu3e, PIONEER, and improved determinations of the electron $g$-2 will be able to definitively test the $X_{17}$ scenario in the near future.
Although we focused on the vector case motivated by the recent PADME analysis, we have also examined scalar, pseudoscalar, and axial-vector couplings, which may be relevant for other $X_{17}$ scenarios or in other contexts. As further data from PADME and related experiments becomes available, the framework explored here will remain a useful tool in scrutinizing the $X_{17}$ interpretation.
Acknowledgments
We thank Simon Knapen and Andre Schöning for useful communications. This work received funding by the INFN Iniziative Specifiche AMPLITUDES and APINE and from the European Union’s Horizon 2020 research and innovation programme under the Marie Sklodowska-Curie grant agreements n. 860881 – HIDDeN, n. 101086085 – ASYMMETRY. This work was also partially supported by the Italian MUR Departments of Excellence grant 2023-2027 “Quantum Frontiers”. The work of LDL and PP is supported by the European Union – Next Generation EU and by the Italian Ministry of University and Research (MUR) via the PRIN 2022 project n. 2022K4B58X – AxionOrigins.